OpenStudy (anonymous):

Write the equation of the line passing through (-3,5) with the slope m=4/3

OpenStudy (lgbasallote):

im just going to write the equation of point slope form dont mind me \(\LARGE (y - y_{1}) = m(x - x_{1})\)

OpenStudy (anonymous):

we have to use the equation of a line which @lgbasallote just stated and plug in your values y1=5 x1=-3, m=4/3 so we get y-5=(4/3)(x+3) and that is just y-5 = (4/3)x +4 and that is just y = (4/3)x -1 :)